OpenStudy (anonymous):

l 1-2x l=9 (l's r absolute value symbols)

OpenStudy (anonymous):



OpenStudy (anonymous):

lets try this \[|1-2x|=9\] \[1-2x=9\] or \[1-2x=-9\] solve those two equations to get your two answers